# Forecasting Influenza Hospitalizations Using a Bayesian Hierarchical Nonlinear Model with Discrepancy

### v1

@@ 6 Conclusion
In this manuscript we introduce a statistical modeling framework which allows for the incorporation of several ILI forecast modeling methods. Specifically, we built upon Osthus et al. Osthus et al. (2019) and introduced a framework for modeling ILI which includes the use of an arbitrary function for modeling the main trajectory of ILI along with modeling the discrepancy. We model flu hospitalizations by incorporating the ILI forecast model into a model forecasting hospitalizations where hospitalization predictions are a linear or quadratic function of ILI.
The simulation study in section 4 suggests the ASG function in ILI modeling may slightly outperform the SIR model according to the LogS and CRPS scoring rules, but in the analysis of the 2023 flu season forecasts, the SIR model was overall superior. The results from both the simulation study and the real data analysis suggest that the addition of a discrepancy component in ILI modeling may improve forecasts especially near the holiday week between Christmas and New Years day. It should not be assumed that these conclusions may be generalized for all locations in the US, weeks of a season, or for future flu seasons.
Forecasting the seasonal influenza outbreak remains a challenging task for forecasters. The general modeling framework in this manuscript is successful under diverse modeling conditions for all locations in the US and may contribute to future forecasting efforts. All forecast models were presented as separate from one another, but in the various locations and times of the flu season, different models perform better than others. To build on the work done here, a natural step forward would be to combine all or a few selected forecasts into an ensemble forecast. Such an ensemble may work to cancel out certain model biases or highlight model strengths, leading to more robust forecasts.
7 Acknowledgments
This work is partially supported by the National Science Foundation under Grant No. 2152117. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
8 Supplementary Material
Supplement to ”Forecasting Influenza Hospitalizations Using a Bayesian Hierarchical Nonlinear Model with Discrepancy”
A. Additional plots for ILI & Hospitalization data for select regions, results from simulation study, and 2023-24 flu analysis
B. Posterior distribution plots for select parameters
C. FluSight forecast competition scoring results

### v2

@@ 6 Conclusion
In this manuscript we introduce a statistical modeling framework which allows for the incorporation of several nonlinear ILI forecast modeling methods. Specifically, we built upon Osthus et al. Osthus et al. (2019) and introduced a framework for modeling ILI which includes the use of an arbitrary function for modeling the main trajectory of ILI along with modeling the discrepancy and specifically incorporate the smooth ASG function. We model flu hospitalizations by incorporating the ILI forecast model into a model forecasting hospitalizations where hospitalization predictions are a linear or quadratic function of ILI. Utlizing the ILI data which has been available for many more years than flu hospitalization, greatly improves one’s ability to forecast hospitalizations, and the borrowing information from all previous seasons through hierarhcical modeling is able to further exploit the ILI data. Including model discrepancy was an important piece of successful forecasting, but it required making thoughtful model constraints and selection of prior distributions.
The simulation study in section 4 suggests the ASGD function in ILI modeling outperforms the SIRD model and out of the box forecast methods according to the WIS scoring rule, and in the analysis of 2023 FluSight forecasts the ASGD_NORM2 modeling scheme outperformed all but one of the competing forecasts when looking at the whole season. Though the ASGD_NORM2 did very well across the whole season, there were a small number of weeks or locations where performance was not as good. It should not be assumed that ASGD_NORM2 forecasts would perform as well in every season given that each flu season will come with its own challenges which the modeling scheme may or may not address.
The models in this manuscript were such that the data in each location were modeled as independent from all other locations. It is reasonable to assume that a disease outbreak in one state will share characterstics with the outbreak in neighboring states, and indeed the behavior of ILI and hospitalizations is more closely related for neighboring states as shown in data plots in the supplementary materials. Notably, the one 2023 FluSight forecast model which outperformed the ASGD_NORM2, the UGA_flucast-INFLAenza, was listed as being a spatial model. And in past seasons, some of the top performing forecast models also explicitly modeled spatial relationships Osthus and Moran (2021). Likewise, the forecasts to which ours was compared were all non-ensemble forecasts which were all outperformed by the ensemble forecasts in FluSight. Futher extensions of the models herein could include explicit modeling of spatial relationships between states and inclusion in an ensemble of other forecast models.